# The Six-Vertex Yang-Baxter Groupoid

### v2

@@ 3. Yang-Baxter equation for the six-vertex model
$u^{\ast}=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{1}^{\ast}(u)&b_{1}^{\ast}(u)&\\ &b_{2}^{\ast}(u)&c_{2}^{\ast}(u)&\\ &&&a_{2}^{\ast}(u)\end{array}\right)=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{2}(u)&-b_{1}(u)&\\ &-b_{2}(u)&c_{1}(u)&\\ &&&a_{2}^{\ast}(u)\end{array}\right),$
@@
$a_{1}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{1}(u)},\qquad a_{1}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{2}(u)}.$
@@ 5. Preparations for the Non-Free-Fermionic Groupoid
Remark 4.
We assume that $u\in\Omega_{b}\cup\Omega_{a}$, leaving the other case to the reader. It may be surpising that a solution always exists, in contrast with Theorem 3.1. But note that $\Delta(u)$ is not defined if $u\in\Omega_{a}$ or $\Omega_{b}$ since the numerator and denominator in (3) and (4) both vanish. Therefore the condition $\Delta(u)=\Delta(v^{\ast})$ has no meaning. This makes it less surprising that the solution exists without condition.
@@ 6. The Six-Vertex Yang-Baxter Groupoid
Since $c_{1}(\mathbf{r})$ anc $c_{2}(\mathbf{r})$ are defined on the entire groupoid. ∎
@@
First suppose $\mathbf{r}\in\Gamma_{b}$, so $b_{1}(r)=b_{2}(r)=0$ and $a_{1}(r)a_{2}(r)=c_{1}(r)c_{2}(r)$. We define the sequence $r_{n}$ so that $a_{1}(r_{n})=a_{1}(r)$ and $a_{2}(r_{n})=a_{2}(r)$. Let $(d_{1},d_{2})=\Delta(\mathbf{r})$. In order that $\Delta(r_{n})=(d_{1},d_{2})$ we need:
@@
Now let us define the groupoid composition $\star$ on $\mathfrak{G}_{\operatorname{nf}}$.
@@
If $\Delta(\mathbf{u})=\Delta(\mathbf{v}^{\ast})$, let $\mathbf{u}\star\mathbf{v}=\mathbf{w}$, where $\mathbf{w}$ is as in Proposition 6.5. In this case we say that $\mathbf{u}\star\mathbf{v}$ is defined
@@
$\begin{array}[]{lll}b_{1}(s\star v)&=&a_{1}(s^{\ast})b_{1}(v)+b_{1}(s)a_{1}(v)\\ &=&(a_{1}(t^{\ast})a_{1}(r^{\ast})-b_{1}(r)b_{2}(t))b_{1}(v)+(b_{1}(r)a_{1}(t)+b_{1}(t)a_{1}(r^{\ast}))a_{1}(v),\\ &&\Delta\\ b_{1}(r\star u)&=&a_{1}(r^{\ast})b_{1}(u)+b_{1}(r)a_{1}(u)\\ &=&a_{1}(r^{\ast})(a_{1}(t^{\ast})b_{1}(v)+b_{1}(t)a_{1}(v))+b_{1}(r)(a_{1}(t)a_{1}(v)-b_{2}(t)b_{1}(v)).\end{array}$
@@
Proposition 6.1 shows that $\mathbf{I}_{d_{1},d_{2}}^{\ast}=\mathbf{I}_{d_{1},d_{2}}$. Therefore $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{I}_{d_{1},d_{2}}$ is defined. It is of the form $\mathbf{i}=(J,D_{1},D_{2})$, where by the uniqueness assertion in Proposition 5.5, $J$ is the unique element of $\overline{\Omega}$ such that $\llbracket I,J,I\rrbracket=0$ is a normalized solution; thus $J=I$, and by Proposition 6.5 we have $(D_{1},D_{2})=\Delta(\mathbf{i})=\Delta(I_{d_{1},d_{2}})=(d_{1},d_{2})$. Therefore $\mathbf{i}=\mathbf{I}_{d_{1},d_{2}}$. We leave the remaining details to the reader. ∎
@@
so $u\star u^{-1}$ is defined and is a constant multiple of $I_{V\otimes V}$. Now $u^{\prime}$ is a constant multiple of $u^{-1}$, so $u\star u^{\prime}$ is a constant multiple of $I_{V\otimes V}$ and it may be checked that $\llbracket u,I,u^{\prime}\rrbracket=0$ is the normalized solution, so indeed $\mathbf{u}\star\mathbf{u}^{\prime}=I_{d_{1},d_{2}}$ where $(d_{1},d_{2})=\Delta(\mathbf{u}^{\prime})=\Delta(\mathbf{u}^{\ast})$. The last assertion follows from associativity and Proposition 6.8.

### v3

@@ 3. Yang-Baxter equation for the six-vertex model
$u^{\ast}=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{1}^{\ast}(u)&b_{1}^{\ast}(u)&\\ &b_{2}^{\ast}(u)&c_{2}^{\ast}(u)&\\ &&&a_{2}^{\ast}(u)\end{array}\right)=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{2}(u)&-b_{1}(u)&\\ &-b_{2}(u)&c_{1}(u)&\\ &&&a_{2}(u^{\ast})\end{array}\right),$
@@
$a_{1}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{1}(u)},\qquad a_{2}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{2}(u)}.$
@@ 5. Preparations for the Non-Free-Fermionic Groupoid
We assume that $u\in\Omega_{b}\cup\Omega_{a}$, leaving the other case (where $v\in\Omega_{b}\cup\Omega_{a}$) to the reader.
@@ 6. The Six-Vertex Yang-Baxter Groupoid
Since $c_{1}(\mathbf{r})$ and $c_{2}(\mathbf{r})$ are defined on the entire groupoid. ∎
@@
First suppose $\mathbf{r}\in\Gamma_{b}$, so $b_{1}(r)=b_{2}(r)=0$ and $a_{1}(r)a_{2}(r)=c_{1}(r)c_{2}(r)$. We will choose the sequence $r_{n}$ so that $a_{1}(r_{n})=a_{1}(r)$ and $a_{2}(r_{n})=a_{2}(r)$. Let $(d_{1},d_{2})=\Delta(\mathbf{r})$. In order that $\Delta(r_{n})=(d_{1},d_{2})$ we need:
@@
Now let us define the groupoid composition $\star$ on $\mathfrak{G}_{\operatorname{nf}}$. If $\Delta(\mathbf{u})=\Delta(\mathbf{v}^{\ast})$, let $\mathbf{u}\star\mathbf{v}=\mathbf{w}$, where $\mathbf{w}$ is as in Proposition 6.5. In this case we say that $\mathbf{u}\star\mathbf{v}$ is defined.
@@
$\begin{array}[]{lll}b_{1}(s\star v)&=&a_{1}(s^{\ast})b_{1}(v)+b_{1}(s)a_{1}(v)\\ &=&(a_{1}(t^{\ast})a_{1}(r^{\ast})-b_{1}(r)b_{2}(t))b_{1}(v)+(b_{1}(r)a_{1}(t)+b_{1}(t)a_{1}(r^{\ast}))a_{1}(v),\\ &&\\ b_{1}(r\star u)&=&a_{1}(r^{\ast})b_{1}(u)+b_{1}(r)a_{1}(u)\\ &=&a_{1}(r^{\ast})(a_{1}(t^{\ast})b_{1}(v)+b_{1}(t)a_{1}(v))+b_{1}(r)(a_{1}(t)a_{1}(v)-b_{2}(t)b_{1}(v)).\end{array}$
@@
Proposition 6.1 shows that $\mathbf{I}_{d_{1},d_{2}}^{\ast}=\mathbf{I}_{d_{1},d_{2}}$. Therefore $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{I}_{d_{1},d_{2}}$ is defined. It is of the form $\mathbf{J}=(J,D_{1},D_{2})$, where by the uniqueness assertion in Proposition 5.5, $J$ is the unique element of $\overline{\Omega}$ such that $\llbracket I,J,I\rrbracket=0$ is a normalized solution; thus $J=I$, and by Proposition 6.5 we have $(D_{1},D_{2})=\Delta(\mathbf{i})=\Delta(I_{d_{1},d_{2}})=(d_{1},d_{2})$. Therefore $\mathbf{J}=\mathbf{I}_{d_{1},d_{2}}$. We leave the remaining details to the reader. ∎
@@
so $u\star u^{-1}$ is defined and is a constant multiple of $I_{V\otimes V}$. Now $u^{\prime}$ is a constant multiple of $u^{-1}$, so $u\star u^{\prime}$ is a constant multiple of $I_{V\otimes V}$ and it may be checked that $\llbracket u,I,u^{\prime}\rrbracket=0$ is the normalized solution, so indeed $\mathbf{u}\star\mathbf{u}^{\prime}=\mathbf{I}_{d_{1},d_{2}}$ where $(d_{1},d_{2})=\Delta(\mathbf{u}^{\prime})=\Delta(\mathbf{u}^{\ast})$. The last assertion follows from associativity and Proposition 6.8.
@@
7. The Five-vertex Groupoid
Let us consider the 5-vertex model in which $b_{2}(v)=0$ for all vertices. Thus $N(v)=a_{1}(v)a_{2}(v)-c_{1}(v)c_{2}(v)$. Assuming that $a_{1}(v)$ and $a_{2}(v)$ are nonzero, we still may define $v^{\ast}$ with
(23) $\begin{array}[]{c}a_{1}(v^{\ast})=\frac{c_{1}(v)c_{2}(v)}{a_{1}(v)},\qquad a_{2}(v^{\ast})=\frac{c_{1}(v)c_{2}(v)}{a_{2}(v)},\\ c_{1}(v^{\ast})=c_{2}(v),\qquad c_{2}(v^{\ast})=c_{1}(v),\qquad b_{1}(v^{\ast})=-b_{1}(v)\end{array}$
and $b_{2}(v^{\ast})=0$. We have
$N(v^{\ast})=-\frac{c_{1}(v)c_{2}(v)}{a_{1}(v)a_{2}(v)}N(v).$
There are parametrized Yang-Baxter equation of five-vertex matrices within the free-fermionic groupoid. But we will ignore thes and instead exhibit a groupoid in the non-free-fermionic part, that is disjoint from the six-vertex groupoid $\mathfrak{G}_{\textrm{nf}}$. Let $\Phi$ be the set of six-vertex matrices such that $c_{1}(v),c_{2}(v),N(v),b_{1}(v),a_{1}(v),a_{2}(v)$ are all nonzero, but $b_{2}(v)=0$. Let $\Phi_{b}$ be the set such that $c_{1}(v),c_{2}(v),a_{1}(v),a_{2}(v)$ are all nonzero but $b_{1}(v)=b_{2}(v)=0$ and $N(v)=a_{1}(v)a_{2}(v)-c_{1}(v)c_{2}(v)=0$. (The set $\Phi_{b}$ coincides with the set $\Omega_{b}$ in Table 1.) Let $\overline{\Phi}=\Phi\cup\Phi_{b}$. In this section we will prove:
Theorem 7.1.
There exists a groupoid $\mathfrak{G}$ with a dominant birational morphism $\pi:\mathfrak{G}\rightarrow\overline{\Phi}$ such that $\llbracket\pi(u),\pi(u\star v),\pi(v)\rrbracket=0$ if $u\star v$ is defined.
Now $\Delta_{1}(v)$ is undefined on the boundary component $\Phi_{b}$. But on $\Phi$ we have
$\Delta_{1}(v)=\frac{N(v)}{a_{1}(v)b_{1}(v)},\qquad\Delta_{1}(v^{\ast})=\frac{N(v)}{a_{2}(v)b_{1}(v)}.$
Proposition 7.2.
Let $u,v\in\Phi$. A necessary and sufficient condition for there to exist $w\in S$ such that $\llbracket u,w,v\rrbracket=0$ is that $\Delta_{1}(u)=\Delta_{1}(v^{\ast})$. If this identity is satisfied, then $w$ is determined up to a constant multiple, $w\in\overline{\Phi}$ and it may be normalized so
(24) $\begin{array}[]{c}c_{1}(w)=c_{1}(u)c_{1}(v),\qquad c_{2}(w)=c_{2}(u)c_{2}(v),\\ a_{1}(w)=a_{1}(u)a_{1}(v),\qquad a_{2}(w)=a_{2}(u)a_{2}(v),\\ b_{1}(w)=b_{1}(u)a_{1}(v)+a_{1}(u^{\ast})b_{1}(v)=a_{2}(u)b_{1}(v)+b_{1}(u)a_{2}(v^{\ast}).\end{array}$
Here the two expressions for $b_{1}(w)$ are equivalent on the assumption that $\Delta_{1}(v)=\Delta_{1}(u^{\ast})$. If $w\in\Phi$ then
$\Delta_{1}(w)=\Delta_{1}(v),\qquad\Delta_{1}(w^{\ast})=\Delta_{1}(u^{\ast}).$
Proof.
This is proved similarly to Theorem 3.1. ∎
Now to produce the groupoid let $\mathfrak{G}=\left\{(v,\varepsilon)\in\overline{\Phi}\times\mathbb{C}^{\times}|N(v)=a_{1}(v)b_{1}(v)\varepsilon\right\}$. Note that the fiber of the projection $\mathfrak{G}\rightarrow\overline{\Phi}$ onto the first component consists of $(v,\Delta_{1}(v))$ over the open set $\Phi$. On the other hand if $v\in\Phi_{b}$ the fiber over $v$ if $v\times\mathbb{C}^{\times}$. By abuse of notation if $v\in\Phi$ then we will identify $v$ with its image $(v,\Delta_{1}(v))\in\mathfrak{G}$.
Let us introduce the notation $\mathbf{v}=(v,\varepsilon)$ for elements of $\mathfrak{G}$, where $v\in\overline{\Phi}$ and $\varepsilon\in\mathbb{C}^{\times}$. Then we extend $\Delta_{1}$ to all of $\mathfrak{G}$ by $\Delta_{1}(\mathbf{v})=\varepsilon$. We extend the map $v\mapsto v^{\ast}$ to all of $\mathfrak{G}$ as follows. If $\mathbf{v}=(v,\Delta_{1}(\mathbf{v}))$ with $v\in\Phi_{b}$ then we define $\mathbf{v}^{\ast}=(v^{\ast},\Delta_{1}(\mathbf{v}^{\ast}))$ where $v^{\ast}$ is given by the formula (23), and
$\Delta_{1}(\mathbf{v}^{\ast})=\frac{a_{1}(v)}{a_{2}(v)}\Delta(\mathbf{v}).$
The composition $\mathbf{u}\star\mathbf{v}$ is defined only if $\Delta_{1}(\mathbf{u})=\Delta_{1}(\mathbf{v}^{\ast})$. In this case we define $\mathbf{u}\star\mathbf{v}=\mathbf{w}$ where $\mathbf{w}=(w,\Delta_{1}(\mathbf{w}))$ with $w$ defined by the equations (24) and $\Delta_{1}(\mathbf{w})=\Delta_{1}(\mathbf{v})$.
The groupoid also has an inverse map, which we will denote $\mathbf{v}\mapsto\mathbf{v}^{\prime}$, in order to avoid confusion with the matrix inverse. If $v\in\Phi$, then $v^{-1},v^{\ast}$ and $v^{\prime}$ are scalar multiples of one another. So $\mathbf{v}^{\prime}$ is not quite the same $\mathbf{v}^{\ast}$ but is a different normalization: if $\mathbf{v}=(v,\varepsilon)$ then $v^{\prime}=\frac{1}{c_{1}(v)c_{2}(v)}v^{\ast}$ and $\mathbf{v}^{\prime}=(v^{\prime},\Delta_{1}(\mathbf{v}^{\ast}))$.
The groupoid $\mathfrak{G}$ is topologized so that $\Delta_{1}$ is continuous. This means that if $\mathbf{v}=(v,\varepsilon)$ a sequence $\{v_{n}\}\subset\Phi$ converges to $\mathbf{v}$ if $v_{n}\rightarrow v$ and $\Delta_{1}(v_{n})\rightarrow\varepsilon$. All of the functions $a_{1},a_{2},b_{1},c_{1},c_{2},N,\Delta_{1},\mathbf{v}^{\ast},\mathbf{v}^{\prime}$ are continuous on $\mathfrak{G}$.
Lemma 7.3.
The composition $\mathbf{u}\star\mathbf{v}$ is defined if and only if $\Delta_{1}(\mathbf{u})=\Delta_{1}(\mathbf{v}^{\ast})$. If this is so, and $\mathbf{w}=\mathbf{u}\star\mathbf{v}$ then $\Delta_{1}(\mathbf{w})=\Delta_{1}(\mathbf{v})$ and $\Delta_{1}(\mathbf{w}^{\ast})=\Delta_{1}(\mathbf{u}^{\ast})$.
Proof.
The first statement is true by definition. For the second statement, these identities are true on the dense open subset $\Phi$ by Theorem 3.1 and the general case follows by continuity. ∎
We note that the operations $\mathbf{v}\mapsto\mathbf{v}^{\ast}$, $\mathbf{v}\mapsto\mathbf{v}^{\prime}$ and $(\mathbf{u},\mathbf{v})\mapsto\mathbf{u}\star\mathbf{v}$ are continuous. To confirm that we have a groupoid we need to prove associativity.
Proposition 7.4.
Let $\mathbf{r},\mathbf{t},\mathbf{u}\in\mathfrak{G}$ such that $\mathbf{r}\star\mathbf{t}$ and $\mathbf{t}\star\mathbf{u}$ are defined. Then $(\mathbf{r}\star\mathbf{t})\star\mathbf{u}$ and $\mathbf{r}\star(\mathbf{t}\star\mathbf{u})$ are both defined, and they are equal.
Proof.
this follows formally from Lemma 7.3 that $(\mathbf{r}\star\mathbf{t})\star\mathbf{u}$ and $\mathbf{r}\star(\mathbf{t}\star\mathbf{u})$ are defined, because $\Delta_{1}(\mathbf{r}\star\mathbf{t})=\Delta_{1}(\mathbf{t})=\Delta_{1}(\mathbf{u}^{\ast})$ and $\Delta(\mathbf{r})=\Delta_{1}(\mathbf{t}^{\ast})=\Delta_{1}((\mathbf{t}\star\mathbf{u})^{\ast})$.
To show that they are equal, we prove this first for $\mathbf{r}=(r,\Delta_{1}(r))$, $\mathbf{t}=(t,\Delta_{1}(t))$ and $\mathbf{u}=(u,\Delta_{1}(u))$ in the interior, so $r,t,u\in\Phi$. We assume that these are in general position so that also $r\star t,t\star u\in\Phi$. To show that $(r\star t)\star u=r\star(t\star u)$, the computation in Proposition 6.6 may be adapted. We have easily
$a_{1}((r\star t)\star u)=a_{1}(r\star t)a_{1}(u)=a_{1}(r)a_{1}(t)a_{1}(u)=a_{1}(r\star(t\star u)),$
and similarly for $a_{2},c_{1}$ and $c_{2}$. For $b_{1}$:
$\begin{array}[]{lll}b_{1}((r\star t)\star v)&=&a_{1}(t^{\ast}\star r^{\ast})b_{1}(v)+b_{1}(r\star t)a_{1}(v)\\ &=&a_{1}(r^{\ast})a_{1}(t^{\ast})b_{1}(v)+(b_{1}(r)a_{1}(t)+a_{1}(r^{\ast})b_{1}(t))a_{1}(v)\\ &=&a_{1}(r^{\ast})(a_{1}(t^{\ast})b_{1}(v)+b_{1}(t)a_{1}(v))+b_{1}(r)a_{1}(t)a_{1}(v)\\ &=&a_{1}(r^{\ast})b_{1}(t\star v)+b_{1}(r)a_{1}(t\star v)\\ &=&b_{1}(r\star(t\star v)).\end{array}$
These calculations show that $(\mathbf{r}\star\mathbf{t})\star\mathbf{u}=\mathbf{r}\star(\mathbf{t}\star\mathbf{u})$ when $\mathbf{r},\mathbf{t}$ and $\mathbf{u}$ are in a dense open set. The general case follows by continuity. ∎
8. Solvable Lattice Models
Assume we have a groupoid parametrized Yang-Baxter equation $\pi:\mathfrak{G}\to\operatorname{End}(V\otimes V)$ with object map $\Delta:\mathfrak{G}\to M$. We will construct solvable lattice models whose Boltzmann weights come from the model. We will show that these models have both row and column solvability. These are familiar results for models associated with group parametrized Yang-Baxter equations. (Compare [4, 10] for example.) However it is worth checking that such standard solvability arguments work for groupoid parametrized Yang-Baxter equations, and can produce models that are not possible using only a subgroup of a given groupoid.
The model will be based on an $m\times n$ grid. At the vertex in the $i$-th row and the $j$-th column vertex we wish to pick an element $\gamma_{ij}$ such that the Boltzmann weights at the vertex are the matrix coefficients of $\pi(\gamma_{ij})$ in the usual way.
$\pi(\gamma_{41})$$\pi(\gamma_{31})$$\pi(\gamma_{21})$$\pi(\gamma_{11})$$\pi(\gamma_{42})$$\pi(\gamma_{32})$$\pi(\gamma_{22})$$\pi(\gamma_{12})$$\pi(\gamma_{43})$$\pi(\gamma_{33})$$\pi(\gamma_{23})$$\pi(\gamma_{13})$$\pi(\gamma_{44})$$\pi(\gamma_{34})$$\pi(\gamma_{24})$$\pi(\gamma_{14})$
Before we show how that $\gamma_{ij}$ can be produced, we consider the properties they need for solvability. Then we will show a systematic procedure fore producing such $\gamma_{ij}$.
We will say that the model is row solvable if for every pair of adjacent rows we may attach an R-matrix $\pi(\rho)$ for some $\rho\in\mathfrak{G}$ such that we can apply the Yang-Baxter equation. Let us illustrate this with a $2\times 2$ grid. After attaching the R-matrix, the grid looks like this:
$\pi(\gamma_{i,1})$$\pi(\gamma_{i,2})$$\pi(\gamma_{i+1,1})$$\pi(\gamma_{i+1,2})$$\pi(\rho)$
After applying the Yang-Baxter equation:
$\pi(\gamma_{i+1,1})$$\pi(\gamma_{i,2})$$\pi(\gamma_{i,1})$$\pi(\gamma_{i+1,2})$$\pi(\rho)$
Here the Yang-Baxter equation is $\llbracket\pi(\rho),\pi(\gamma_{i,1}),\pi(\gamma_{i+1,1})\rrbracket=0$, so we need $\rho\star\gamma_{i+1,1}=\gamma_{i,1}$ in $\mathfrak{G}$, or $\rho=\gamma_{i,1}\gamma_{i+1,1}^{\prime}$ in $\mathfrak{G}$. Then But we want to use the Yang-Baxter equation again so we need $\rho=\gamma_{i,2}\star\gamma_{i+1,2}^{\prime}$. Thus we need
$\gamma_{i,1}\star\gamma_{i+1,1}^{\prime}=\gamma_{i,2}\star\gamma_{i+1,2}^{\prime}$
in $\mathfrak{G}$. In order to finish the train argument (running the R-matrix past every column) we therefore need:
Desideratum 8.1 (Requirement for row solvability).
For every $i,i+1$ we need $\gamma_{i,j}\star\gamma_{i+1,j}^{\prime}$ to be defined and independent of $j$.
Column solvability would be handled the same way, attaching an R-matrix above the $j,j+1$ columns. Similarly:
Desideratum 8.2 (Requirement for column solvability).
For every $j,j+1$ we need $\gamma_{i,j}^{\prime}\star\gamma_{i,j+1}$ to be defined and independent of $i$.
Proposition 8.3.
Row solvability is equivalent to column solvability.
Proof.
Assume row solvability, so $\gamma_{i,j}\star\gamma_{i+1,j}^{\prime}$ is defined and $\gamma_{i,j}\star\gamma^{\prime}_{i+1,j}=\gamma_{i,j+1}\star\gamma_{i+1,j+1}^{\prime}$. By the inverse groupoid axiom $\gamma_{i,j}^{\prime}\star(\gamma_{i,j}\star\gamma_{i+1,j}^{\prime})$ is defined and equals $\gamma_{i+1,j}^{\prime}$. Therefore $\gamma_{i+1,j}^{\prime}=\gamma_{i,j}^{\prime}\star\gamma_{i,j+1}\star\gamma_{i+1,j+1}^{\prime}$. Another application of the inverse axiom gives $\gamma_{i+1,j}^{\prime}\star\gamma_{i+1,j+1}=\gamma_{i,j}^{\prime}\star\gamma_{i+1,j}$. ∎
Now let us show how to construct models satisfying these desiderata. We fix an element $d\in\mathfrak{G}$, and choose $\phi_{1},\cdots,\phi_{n}\in\mathfrak{G}$ such that $\Delta(\phi_{i})=d$, and $\psi_{1},\cdots,\psi_{m}$ such that $\Delta(\psi_{j}^{\prime})=d$. Then $\Delta(\phi_{i})=\Delta(\psi_{j}^{\prime})$ so $\gamma_{i,j}:=\phi_{i}\star\psi_{j}$ is defined for all $i,j$.
Proposition 8.4.
With this construction $\gamma_{i,j}\star\gamma_{i+1,j}^{\prime}$ is defined and independent of $j$. Hence the row and column solvability desiderata are satisfied and the model is solvable.
Proof.
Let us show that $\gamma_{i,j}\star\gamma_{i+1,j}^{\prime}$ is defined and equals $\phi_{i}\star\phi_{i+1}^{\prime}$. Indeed, $\Delta(\gamma_{i,j})=\Delta(\phi_{i}\star\psi_{j})=\Delta(\psi_{j})=\Delta(\gamma_{i+1,j})$. This implies that $\gamma_{i,j}\star\gamma_{i+1,j}^{\prime}$ is defined. It equals
$\phi_{i}\star\psi_{j}\star(\phi_{i+1}\star\gamma_{j})^{\prime}=\phi_{i}\star\psi_{j}\star\psi_{j}^{\prime}\star\phi_{i+1}^{\prime}=\phi_{i}\star\phi_{i+1}^{\prime}.$
This is independent of $j$ proving row solvability. ∎
We have therefore constructed a class of solvable lattice models using a groupoid parametrized Yang-Baxter equation. We will now address the question of whether these are truly more general than models that could be constructed using group parametrized Yang-Baxter equations. After all, we have imposed a condition that $\Delta(\phi_{i})=\Delta(\psi_{j}^{\prime})=d$ for some fixed element $d\in\mathfrak{G}$. How strongly does this constrain the $\gamma_{i,j}$?
Before addressing this question, we mention a criterion for a groupoid to be a disjoint union of groups. If $\mathfrak{G}$ is a disjoint union of groups, and $v\in\mathfrak{G}$ then $v\star v$ is always defined, and so $\Delta(v)=\Delta(v^{\prime})$. From this we see that if $\llbracket u,v,w\rrbracket=0$ we have $\Delta(u)=\Delta(v)=\Delta(w)$.
On the other hand, in the six- and five-vertex groupoids, $\Delta(v)$ and $\Delta(v^{\prime})$ are decoupled. That is, if we define $M_{d}=\{\Delta(v^{\prime})|v\in M,\Delta(v)=d\}$, the set $M_{d}$ is large.
Now consider that
$\Delta(\gamma_{ij})=\Delta(\psi_{j}),\qquad\Delta(\gamma_{ij}^{\prime})=\Delta(\phi_{i}^{\prime}).$
Although $\Delta(\phi_{i})=\Delta(\psi_{j}^{\prime})=d$ for all $i,j$, the values of $\Delta(\psi_{j})$ and $\Delta(\phi_{i}^{\prime})$ can take arbitrary values from $M_{d}$. Thus $\Delta(\gamma_{ij})$ and $\Delta(\gamma_{ij}^{\prime})$ can also be made arbitrary elements of $M_{d}$, subject only to the constraint that $\Delta(\gamma_{ij})$ depends only on the column $j$, while $\Delta(\gamma_{ij}^{\prime})$ depends only on the row $i$. This is in contrast with classical models such as those in [1, Chapter 9], [10] or [4], where $\Delta(\gamma_{ij})$ and $\Delta(\gamma_{ij}^{\prime})$ are necessarily the same for all $i,j$.